OpenStudy (anonymous):

Electron exhibiting a particle wavelength of 3.35 x 10-9 are emitted from hydrogen atoms after interaction with high energy photons. Assuming that the electrons in the hydrogen atom were in their ground state prior to interaction, determine the wavelength of the photon which generated the free electrons. If someone knows the formula i'm supposed to use, i'd appreciate it.

OpenStudy (anonymous):

E = h*c/L where E is the energy, h is planck's constant, c is the speed of light, and L is the wavelength. Use h = 6.626*10^-34, c = 3*10^8, and put you L in units of meters. so 3.3 um = 3.3*10^-6 and .155 nm = .155*10^-9

OpenStudy (chmvijay):

determine the wavelength of the photon which generated the free electrons.? is the question

OpenStudy (anonymous):

9.03466211997e-08 is the answer according to the solutions

OpenStudy (anonymous):

Thats about right (:

OpenStudy (chmvijay):

http://openstudy.com/study#/updates/525ad22ee4b002bdb08fd300 please find it here

OpenStudy (anonymous):

Kay thanks, real helpful.